OpenStudy (selena2345):

Andre made a cleaning solution using 7/8 cup of water and 1/4 cup of vinegar.He then used 1/3 cup of the solution to wash the windows in his home. Write an expression showing how to find the amount of cleaning solution, in cups that Andre has left. Evaluate the expression you wrote.

OpenStudy (anonymous):

it would be \[\frac{7}{8}+\frac{1}{4}-\frac{1}{3}\]